dx Evaluate by using the limit definition of the definite integral

\int_{1}^{7}(6-2x)dx

Evaluate by using the limit definition of the definite integral
Answer #1

given thata [a b] = [2,7 An 7-1 = 6 h I + P = f(G) = 6-2(1444) = (4 *) 7. E lim S (6-24 ) che (4- 12,2 x 4 hooo 7 lim [24 72